# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ 6 Empirical Results
Table 1: We report the (test_accuracy, fairness_violation) for evaluating the performance on the CelebA dataset with four binary targets, including Smiling, Attractive, Young, and Big_Nose. We select gender as the sensitive attribute.
$\epsilon=0.05$ Smiling Attractive
DP EOp DP EOp
Base (ERM) (0.789 ± 0.003, 0.127 ± 0.007) (0.789 ± 0.003, 0.067 ± 0.003) (0.664 ± 0.005, 0.251 ± 0.022) (0.658 ± 0.009, 0.179 ± 0.014)
Random (0.837 ± 0.017, 0.133 ± 0.001) (0.840 ± 0.020, 0.059 ± 0.005) (0.693 ± 0.002, 0.372 ± 0.006) (0.693 ± 0.012, 0.250 ± 0.011)
BALD (0.856 ± 0.024, 0.140 ± 0.012) (0.854 ± 0.025, 0.062 ± 0.005) (0.712 ± 0.015, 0.431 ± 0.020) (0.712 ± 0.016, 0.297 ± 0.008)
ISAL (0.867 ± 0.022, 0.138 ± 0.010) (0.867 ± 0.022, 0.057 ± 0.007) (0.713 ± 0.009, 0.410 ± 0.003) (0.713 ± 0.009, 0.275 ± 0.011)
JTT-20 (0.698 ± 0.018, 0.077 ± 0.006) (0.679 ± 0.013, 0.045 ± 0.002) (0.621 ± 0.015, 0.094 ± 0.004) (0.605 ± 0.013, 0.083 ± 0.011)
FIS (0.848 ± 0.025, 0.084 ± 0.055) (0.845 ± 0.034, 0.032 ± 0.007) (0.644 ± 0.025, 0.200 ± 0.060) (0.674 ± 0.015, 0.112 ± 0.025)
$\epsilon=0.05$ Young Big_Nose
DP EOp DP EOp
Base(ERM) (0.755 ± 0.002, 0.190 ± 0.017) (0.759 ± 0.005, 0.102 ± 0.005) (0.752 ± 0.024, 0.198 ± 0.034) (0.755 ± 0.022, 0.206 ± 0.018)
Random (0.763 ± 0.008, 0.158 ± 0.016) (0.698 ± 0.109, 0.075 ± 0.021) (0.760 ± 0.009, 0.177 ± 0.014) (0.757 ± 0.004, 0.190 ± 0.029)
BALD (0.776 ± 0.021, 0.165 ± 0.019) (0.775 ± 0.020, 0.076 ± 0.007) (0.777 ± 0.004, 0.184 ± 0.016) (0.765 ± 0.003, 0.209 ± 0.014)
ISAL (0.781 ± 0.020, 0.180 ± 0.014) (0.781 ± 0.020, 0.084 ± 0.006) (0.782 ± 0.001, 0.148 ± 0.059) (0.782 ± 0.001, 0.154 ± 0.080)
JTT-20 (0.774 ± 0.026, 0.167 ± 0.016) (0.774 ± 0.024, 0.083 ± 0.007) (0.771 ± 0.014, 0.191 ± 0.036) (0.758 ± 0.026, 0.223 ± 0.018)
FIS (0.763 ± 0.004, 0.104 ± 0.059) (0.773 ± 0.003, 0.041 ± 0.015) (0.779 ± 0.009, 0.089 ± 0.076) (0.780 ± 0.013, 0.046 ± 0.072)
@@ 6.1 Experimental Setup
We evaluate the performance of our algorithm on four real-world datasets across three different modalities: CelebA (Liu et al., 2015), UCI Adult (Asuncion and Newman, 2007) and COMPAS (Angwin et al., 2016). We present more details about the datasets in Appendix D.1. We implement the fairness constraint $\phi(\cdot)$ using three common group fairness metrics: difference of demographic parity (DP), difference of equality of opportunity (EOp) and difference of equal odds (EOd).
We compare our method with five baselines: 1) Base (ERM): directly train the model solely on the labeled dataset $P$; 2) Random: train the model on the dataset $P$ and a subset of $Q$ by random sampling; 3) BALD (Branchaud-Charron et al., 2021): active sampling according to the mutual information; 4) ISAL (Liu et al., 2021a): selects unlabeled examples based on the calculated influence in an active learning setting. We apply model predictions as pseudo-labels. 5) Just Train Twice (JTT) (Liu et al., 2021b): reweighting those misclassified examples for retraining. Here, we examine a weight of 20 for misclassified examples, marked as JTT-20.
Recall that we present the average result of the classifier $\mathbf{w}_{t}$ outputs from Algorithm 1. Additional details on datasets and hyper-parameters are in Appendix D.
Table 2: The performance results of (test_accuracy, fairness_violation) on the Adult dataset. The selected sensitive attribute is age.
$\epsilon=0.05$ Income
DP EOp
Base(ERM) (0.665 ± 0.045, 0.255 ± 0.041) (0.665 ± 0.045, 0.115 ± 0.036)
Random (0.765 ± 0.021, 0.209 ± 0.042) (0.758 ± 0.027, 0.127 ± 0.013)
BALD (0.767 ± 0.019, 0.203 ± 0.017) (0.703 ± 0.111, 0.117 ± 0.013)
ISAL (0.765 ± 0.020, 0.215 ± 0.011) (0.755 ± 0.028, 0.128 ± 0.013)
JTT-20 (0.751 ± 0.013, 0.262 ± 0.020) (0.742 ± 0.018, 0.149 ± 0.021)
FIS (0.766 ± 0.013, 0.214 ± 0.009) (0.757 ± 0.034, 0.113 ± 0.017)
$\epsilon=0.05$ Compas-Recidivism
DP EOp
Base (ERM) (0.675 ± 0.005, 0.333 ± 0.008) (0.675 ± 0.005, 0.267 ± 0.010)
Random (0.689 ± 0.007, 0.305 ± 0.023) (0.686 ± 0.016, 0.253 ± 0.035)
BALD (0.688 ± 0.011, 0.313 ± 0.012) (0.686 ± 0.015, 0.256 ± 0.031)
ISAL (0.697 ± 0.002, 0.308 ± 0.025) (0.698 ± 0.004, 0.274 ± 0.022)
JTT-20 (0.646 ± 0.009, 0.240 ± 0.016) (0.630 ± 0.024, 0.141 ± 0.028)
FIS (0.690 ± 0.002, 0.299 ± 0.029) (0.694 ± 0.002, 0.241 ± 0.035)
6.2 Performance Results
@@
Note that all the experimental results presented subsequently are from three separate trials, each conducted with distinct random seeds. Due to space limits, we provide a full version of the experimental results (tables) in the Appendix D.
Figure 1: The impact of label budgets on the test accuracy & DP gap in the CelebA and Adult dataset, respectively. The binary classification target is Smiling for the CelebA dataset.
@@
Results on image datasets. Initially, we train a vision transformer using a patch size of $(8,8)$ on the CelebA face attribute dataset (Liu et al., 2015). We select four binary classification targets, including Smiling, Attractive, Young, and Big Nose. The sensitive attribute is gender. 2% of the labeled data is allocated for training, while the remaining 98% is reserved for sampling purposes. Then, 10% of the test data is randomly set aside as a hold-out validation set. For ease of computation, only the last two layers of the model are used to calculate the influence of prediction and fairness components. Table 1 report the obtained accuracy and the corresponding values of fairness metrics. One main observation is that our proposed method FIS outperforms baselines with a significant margin on three fairness metrics, while maintaining the same accuracy level. This improvement, as indicated by our findings in Theorem 4.2, can be attributed to FIS assigning priority to new examples based on the influence of fairness, thus averting potential accuracy reduction by assessing the corresponding influence of prediction.
@@
Results on tabular datasets. Next, we work with multi-layer perceptron (MLP) with two layers trained on the Adult dataset (Asuncion and Newman, 2007) and Compas dataset (Angwin et al., 2016), respectively. We select age as the sensitive attribute for the Adult dataset and race for the COMPAS dataset. For both of the datasets, we resample the data to balance the class and group membership (Chawla et al., 2002). We randomly split the whole dataset into a training and a test set in a ratio of 80 to 20. Then, we randomly re-select 20% of the training set for initial training and the remaining 80% for sampling. Also, 20% examples of the test set are selected to form a validation set. The MLP model is a two-layer ReLU network with a hidden size of 64. We utilize the whole model parameters to compute the influence of prediction and fairness for examples. Table 2 and Table 3 summarize the key results of the Adult and Compas datasets, respectively. On Adult dataset, we observe that our sampling method achieves the lowest violation for equality of opportunity and has a comparable performance for the demographic parity metric. Besides, our algorithm achieves a much better accuracy-fairness trade-off than the other compared methods on the COMPAS dataset. Though the JTT-20 method indeed has a lower fairness violation, we highlight that they suffer from a dramastic accuracy drop than the other baselines.
6.3 Impact of Label Budgets
@@
We conduct an ablation study to examine how the varying label budgets $r$ affects the trade-off between accuracy and fairness. For ease of comparison, we adhere to a consistent label budget per round to illustrate their respective impacts. As shown in Figure 1, our method consistently preserves a lower fairness violation than the BALD and ISAL baselines with a similar test accuracy. While we observe that the JTT-20 algorithm is able to achieve a near-zero fairness violation under a limited budget on the CelebA dataset, we argue that the prediction accuracy is rather uninformative (about 50%). Conversely, on the Adult dataset, both accuracy and fairness violation converge to similar numerical values when the budget is lower than 20, suggesting a potential overfitting of the model to insufficient training examples. With a larger budget, our algorithms outperform other baseline methods, achieving higher accuracy and lower demographic disparity. We defer more experiment details in the Appendix D.3.
7 Concluding Remarks
In this work, we explore the training of fair classifiers without imposing fairness constraints, aiming to prevent the potential exposure of sensitive attributes. Our theoretical finding affirms that by using traditional training on suitably shifted dataset distributions, we can decrease the bounds of fairness disparity and model generalization error simultaneously. Motivated by the insights from our results, we propose a fair influential sampling method FIS to inquiry data samples from a large unlabeled dataset to progressively shift the original training data during training. Importantly, the sensitive attribute of new examples is neither accessed during sampling nor used in the training phase. The privilege of our method is that we acquire the data examples that may cause negative impacts on the fairness violations during the sampling process. Excluding these specific data examples during training is likely to contribute to mitigating unfairness at test time. Empirical experiments on real-world data validate the efficacy of our proposed method.
@@ 7 Concluding Remarks
Nonetheless, we recognize that our method has limitations. Although the proposed sampling algorithm does not require the sensitive attribute information from the massive unlabeled data, it relies on a clean and informative validation set which contains the sensitive attributes of data examples. We consider this as a reasonable requirement in practice, given the relatively modest size of the validation set.
@@ Appendix A More Details of Related Work
Active learning. The core idea of active learning is to rank unlabeled instances by developing specific significant measures, including uncertainty [Lewis and Catlett, 1994, Lindenbaum et al., 2004], representativeness [Dasgupta and Hsu, 2008], inconsistency [Wang et al., 2012], variance [Hoi et al., 2006], and error [Roy and McCallum, 2001]. Each of these measures has its own criterion to determine the importance of instances for enhancing classifier performance. For example, uncertainty considers the most important unlabeled instance to be the nearest one to the current classification boundary.
@@
A related line of work [Liu et al., 2021a, Wang et al., 2022b, Hammoudeh and Lowd, 2022, Xie et al., 2023] concentrates on ranking unlabeled instances based on the influence function. Compared to these studies with a focus on the prediction accuracy, our work poses a distinct challenge taking into account fairness violations. We note that adopting a particular sampling strategy can lead to distribution shifts between the training and testing data. What’s worse, even though fairness is satisfied within training dataset, the model may still exhibit unfair treatments on test dataset due to the distribution shift. Therefore, it becomes imperative for the sampling approach to also account for its potential impacts on fairness.

### v3

@@ 6.1 Experimental setup
We evaluate the performance of our algorithm on three real-world datasets across three different modalities: CelebA [49], UCI Adult [8] and Compas [6]. We implement the fairness loss $\phi(\cdot)$ based on three common group fairness metrics: difference of demographic parity (DP), difference of equality of opportunity (EOp), and difference of equal odds (EOd). We compare our method with five baselines: 1) Base (ERM): directly train the model on the training dataset $P$; 2) Random: train the model on dataset $P$ and randomly sampled data from $U$ with inquired true labels; 3) BALD [12]: active sampling according to the mutual information; 4) ISAL [48]: selects unlabeled examples based on the calculated influence in an active learning setting. We apply model predictions as pseudo-labels; 5) Just Train Twice (JTT) [47]: reweighting those misclassified examples for re-training. Here, we examine a weight of 20 for misclassified examples, marked as JTT-20. Recall that we present the average result of the classifier $\mathbf{w}_{t}$ outputs from Algorithm 1. The general term “fairness violation” is utilized to quantify the absolute differences based on fairness metrics, such as DP and EOd. More details on datasets and hyper-parameters are provided in Appendix E.1.
6.2 Main results
@@
Note that all the experimental results presented subsequently are from three independent trials, each conducted with distinct random seeds. We present the primary results as tuples in the form (test_accuracy, fairness_violation) to facilitate comparison of the fairness-accuracy tradeoff. Due to space limits, we provide a full version of the experimental results (tables) in Appendix E.2.
Results on image datasets
@@
Initially, we train a vision transformer using a patch size of $(8,8)$ on the CelebA face attribute dataset [49]. We select four binary classification targets, including Smiling, Attractive, Young, and Big Nose. The sensitive attribute is gender. 2% of the labeled data is allocated for training, while the remaining 98% is reserved for sampling purposes. Then, the test dataset is split into two independent portions: a new test set and a validation set, with 10% of the test data randomly designated as the hold-out validation set. For ease of computation, only the last two layers of the model are used to calculate the influence of accuracy and fairness components. For Figure 2, one main observation is that FIS outperforms baselines with a significant margin on three fairness metrics while maintaining the same accuracy level. This improvement, as indicated in Theorem 5.2, can be attributed to FIS assigning priority to new examples based on the fairness influence, then avoiding accuracy reduction via their accuracy influence.
Figure 2: Main results on CelebA, Adult and Compas datasets. The Y axis shows fairness_violation; X axis denotes test_accuracy. CelebA: Four binary targets: Smiling, Attractive, Young, and Big_Nose; Sensitive attribute: gender. Adult: Binary target: Income; Sensitive attribute: Age. Compas: Binary target: Recidivism; Sensitive attribute: Race. We select two fairness metrics DP and Eop to measure fairness violations for each setting. The vertical dotted line at the random baseline accuracy helps easily identify which results achieve fairness without sacrificing performance (accuracy).
Results on tabular datasets
@@
Next, we work with multi-layer perceptron (MLP) with two layers trained on the Adult [8] and Compas dataset [6], respectively. We select age as the sensitive attribute for the Adult dataset and race for the Compas dataset. For two datasets, we resample the data to balance the class and group membership [17]. The whole dataset is split into training and test sets at a 4:1 ratio. Then, we randomly re-select 20% of the training set for initial training and the remaining 80% for sampling. Also, 20% examples of the test set are selected to form a validation set. The MLP model is a two-layer ReLU network with a hidden size of 64. We utilize the whole model parameters to compute the influence of accuracy and fairness for examples. Figure 2 summarizes the main results of the Adult and Compas datasets. On the Adult dataset, we observe that our sampling method achieves the lowest violation for equality of opportunity and has a comparable performance for the DP metric. Besides, our algorithm achieves a much better accuracy-fairness trade-off than other baselines on the Compas dataset. JTT-20 achieves a lower fairness violation with the price of a significant accuracy drop compared to other baselines.
6.3 Ablation study
What is the impact of label budgets?
@@ 6.3 Ablation study
Here, we examine how the varying label budgets $r$ affect the trade-off between accuracy and fairness. For ease of comparison, we adhere to a consistent label budget per round to illustrate their respective impacts. As shown in Figure 3, our method consistently preserves a lower fairness violation than the BALD and ISAL baselines with a similar test accuracy. While we observe that the JTT-20 algorithm can achieve a near-zero fairness violation under a limited budget on the CelebA dataset, we argue that the model accuracy is rather uninformative (about 50%). More empirical results can be found in Appendix E.3.
How does the validation set size affect the performance?
We further explore the impact of adjusting the validation set size on our algorithm’s performance. We present the test accuracy and fairness violations across different validation set sizes on the CelebA dataset. Note that the default validation set size is set to 1% of the whole dataset size. In particular, the minimum scale of the validation set size is set to $\nicefrac{{1}}{{5}}\times$ (nearly 400 CelebA images). The results in Figure 3 indicate that our algorithm still retains the test accuracy and fairness violation when we vary the validation set size. Additional results conducted on Adult and Compas datasets are provided in Appendix E.4.
CelebA - Smiling
[…]
1×1$\times$1 × (0.848, 0.084) (0.876, 0.031) (0.864, 0.030)
$\nicefrac{{1}}{{2}}\times$ (0.872, 0.105) (0.891, 0.042) (0.880, 0.028)
$\nicefrac{{1}}{{5}}\times$ (0.872, 0.117) (0.863, 0.057) (0.886, 0.028)
Figure 3: Left: The impact of label budgets on the test accuracy & DP gap in the CelebA dataset. Right: The impact of the validation set size on (test_accuracy, fairness_violation) results.
7 Conclusions and limitations
In this work, we are interested in facilitating ML models that mitigate group fairness disparity without harming model accuracy. To achieve this, different from current active sampling methods, we propose a tractable fair influential sampling method FIS, which avoids the need for training group annotations during the sampling or training phase, thereby preventing the potential exposure of sensitive information. In particular, this algorithm acquires data samples from a large dataset for training based on the influence of fairness and accuracy evaluated on the auxiliary validation dataset. Empirical experiments on real-world data validate the efficacy of our proposed method.
@@ 7 Conclusions and limitations
Nonetheless, we recognize that our method has limitations. Although the proposed sampling algorithm does not require sensitive attribute information from the massive data, it relies on a clean and informative validation set that contains the sensitive attributes of data examples. We consider this as a reasonable requirement in practice, given the relatively modest size of the validation set. Besides, one potential concern is that the sampling strategy may become inefficient when the collected validation set is noisy. However, this practical issue can be heavily alleviated by using loss correction methods [54, 84] or noise-tolerant fairness loss functions to rectify the error terms [53, 73, 29]. In future work, we aim to address this limitation by developing more robust sampling strategies that can perform effectively even with noisy validation sets.
Acknowledgment
This work is partially supported by the National Science Foundation (NSF) under grants IIS-2143895, IIS-2040800, IIS-2416896, and CCF-2023495. Additionally, Pang and Qian were partially supported by NSF Grants 2322919, 2420632, 2426031, and 2114113.
Broader Impact
This paper presents work whose goal is to advance the field of fairness in machine learning. There are many potential societal consequences of our work. While the proposed algorithm does intend to infer sensitive attributes of data examples that may be protected by privacy regulations, it does not necessitate direct access to such sensitive information. On the other hand, our work can serve as an effective approach leading to mitigating the disparity with a limited annotation budget. We have thoroughly examined the potential ethical implications of our work and, based on our assessment, do not identify any issues that we deem necessary to emphasize here specifically.
@@ Acknowledgment
The Appendix is organized as follows.
  • $\bullet$
    Section A provides more details of the related work.
  • $\bullet$
    Section B explores the relationship between our proposed fairness notion risk disparity and common fairness metrics, such as DP and EOd. In particular, we provide the full proof for Proposition 3.1.
  • $\bullet$
    Section C provides a detailed analysis of the FIS algorithm including 1) evaluating first-order influence estimations against real influence, 2) a comparative analysis of computational costs, and 3) the exploration of the labeling strategies.
  • $\bullet$
    Section D presents the full proofs for the Lemmas and Theorems shown in Section 4 and Section 5.
  • $\bullet$
    Section E presents detailed descriptions of all datasets, corresponding parameter settings, and full version of the experimental results. In particular, to demonstrate FIS’s advantage at the same levels of information, we introduce a new baseline called Random+Val.
  • $\bullet$
    Section F discusses the privacy concern potentially caused by the correlations between non-demographic and demographic features.
@@ Appendix A More details of related work
Active learning The core idea of active learning is to rank unlabeled instances by developing specific significant measures, including uncertainty [41, 45], representativeness [22], inconsistency [70], variance [34], and error [59]. Each of these measures has its criterion to determine the importance of instances for enhancing classifier performance. For example, uncertainty considers the most important unlabeled instance to be the nearest one to the current classification boundary. A related line of work [48, 71, 31, 76] concentrates on ranking unlabeled instances based on the influence function. Compared to these studies with a focus on prediction accuracy, our work poses a distinct challenge taking into account fairness violations. We note that adopting a particular sampling strategy can lead to distribution shifts between the training and testing data. What’s worse, even though fairness is satisfied within the training dataset, the model may still exhibit unfair treatments on the test dataset due to the distribution shift. Therefore, it becomes imperative for the sampling approach to also account for its potential impacts on fairness.
Pareto optimality In the field of fairness in machine learning, Pareto optimality indicates the theoretical frontier of fairness accuracy tradeoff, meaning that fairness can not be improved without worsening model accuracy. Existing methods primarily focus on seeking the Pareto frontier of fairness accuracy tradeoff for the neural network classifier, instead of reaching a better one. For example, Balashankar et al. [10] first analyzes the Pareto optimality for classifiers within the context of fairness constraints. Wang et al. [72] explores the multi-dimensional Pareto frontiers of the fairness-accuracy tradeoff in the multi-task setting. Another works [51, 50] target to obtain a Pareto efficient classifier to reduce worst-case group risks by formulating group fairness as a multiple-objective optimization problem, where each group risk is an objective function.